OpenStudy (anonymous):

plz help me Solve for x. 4x + 6 = 38 A. 3.5 B. 8 C. 11 D. 128

pooja195 (pooja195):

Any clue on what the first step might be?

OpenStudy (anonymous):

no i dont think so because when i solved it i got the wrong answer

pooja195 (pooja195):

Can you show me what you did? And what you got?

OpenStudy (anonymous):

i got 3.8 :(

pooja195 (pooja195):

Hmm ok 4x + 6 = 38 First step subtract both sides by 6 6-6=0 so we are left with 4x 38-6=?

OpenStudy (anonymous):

ohhhhhhh

pooja195 (pooja195):

Solve from there tell me what u get

OpenStudy (anonymous):

\[4x=32 \] \[32\div4=8\] \[x=8\] is that right

pooja195 (pooja195):

Well done and thanks for showing your steps! :D

OpenStudy (anonymous):

np and thank youuuu

pooja195 (pooja195):

yw ^_^